OpenStudy (anonymous):

An engine burns 1.4 kg of gasoline per hour and yields 3.5 kW of power. Calculate the efficiency of the engine when the heating value of gasoline is 43 MJ/kg.

OpenStudy (anonymous):

i dont even know where to start

OpenStudy (anonymous):

ANY help, even a formula or a rule or a hint, anything will help

OpenStudy (anonymous):

i just know efficiency = input/output

OpenStudy (anonymous):

Yep, calm down.,,

OpenStudy (anonymous):

oh, sorry, it is output/input

OpenStudy (anonymous):

ok,.... i am calm

OpenStudy (anonymous):

The heating value is the amount of heat energy produced by complete combustion in a given quantity. Take a basis of 1 hour. Then your engine burns 1.4kg of gasoline, which yields a maximal amount of thermal energy to use for work as\[43MJ/kg \times 1.4 kg=60.2MJ\]In one hour, your engine produces from this energy, \[\frac{3500J}{s}\times \frac{3600s}{h}=12.6MJ\]The efficiency is the amount of work per unit energy input, which is \[\eta = \frac{12.6MJ}{60.2MJ}=0.209\]to 3 significant figures.

OpenStudy (anonymous):

This is the first time I've used 'heating value' for anything like this, so if the answer's wrong, I couldn't find an appropriate definition.

OpenStudy (anonymous):

so, can i say that efficiancy is 20.9%?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok, thank you, i already faned you long time ago, so cant fan again :(

OpenStudy (anonymous):

Yeah I know - no worries :)

OpenStudy (anonymous):

The answer makes sense - I've seen values between 10 and 50% for internal combustion engines.

OpenStudy (anonymous):

then it is probably correct, i also know that cobustion engines are very unefficient

OpenStudy (anonymous):

hey anwar

OpenStudy (anonymous):

hi :)

OpenStudy (anonymous):

want a fan?

OpenStudy (anonymous):

would you be one if I say yes?

OpenStudy (anonymous):

...

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

you got it

OpenStudy (anonymous):

thanks!!

OpenStudy (anonymous):

anwar, can you please fan lokisan, because he helped me alot and i can only fan him once

OpenStudy (anonymous):

lol you really care about others :P.. I already did anyways

OpenStudy (anonymous):

so you faned me just as an exchange for me to fan the guy huh? :(

OpenStudy (anonymous):

nah, would have faned you anyway

OpenStudy (anonymous):

lol andy, you and your fanning job~ ^_^